OpenStudy (anonymous):

Find the point on the terminal side of θ = that has an x coordinate of -1. How?! HELP

OpenStudy (anonymous):

@tcarroll010 Could you help this is Pre-Calculus?

OpenStudy (jdoe0001):

θ = that ?

OpenStudy (jdoe0001):

you kinda got a blank there

OpenStudy (anonymous):

sorry\[-3\pi/4\]

OpenStudy (anonymous):

I'm using the Unit Circle but I'm not sure how to get the pint.

OpenStudy (anonymous):

Would the terminal side be 135 degrees?

OpenStudy (jdoe0001):

well, negative values go COUNTERCLOCKWISE so they don't go upward and to the left negative angles are measured downwards and to the left for example \(\cfrac{3\pi}{2} = -\cfrac{\pi}{2}\)

OpenStudy (jdoe0001):

wait, I said that backwards, they go CLOCKWISE

OpenStudy (jdoe0001):

or the same way a CLOCK hands move

OpenStudy (anonymous):

So −3π/4 = pi/4

OpenStudy (jdoe0001):

well, 4/4 of pi, is a "pi", so 3/4 of pi, will be less than that by 1/4 so 

OpenStudy (anonymous):

Ok I understand that "well, 4/4 of pi, is a "pi", so 3/4 of pi, will be less than that by 1/4" but what is the image showing? One piece of pi is 4/4 which you put but